Grade 11Mechanics

A projectile is thrown with a velocity of 20m/s at an angle of 60° with the horizontal .after how much time the velocity vector will make an angle 45° with the horizontal.

Shiba prasad das
9 Years agoGrade 11

4 Answers

DHRUV NARAYAN TIWARI
9 Years ago
A projectile is thrown with a velocity of 20m/s at an angle of 60° with the horizontal .after 5-6 seconds the velocity vector will make an angle 45° with the horizontal.
 
kishan
8 Years ago
Sala e kaun chutia h Jo answer diya h 5 se 6 second me ho jayega sala reason kaun dega stupid...................
Azeez
8 Years ago
Haha lol 5-6 seconds great answer i appreciate ur effort towards the ans 😂😂 keep it up!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
Azeez
8 Years ago
For 45 degree we will use tan 45=Vertical velocity÷horizontal velocityVertical velocity=20 sin60Horizontal velocity=20 cos60ThereforeVf=Vi-gt20 cos 60=20 sin60-gt10=10root-10tt=root3-1